OpenStudy (anonymous):

A hand consists of 5 cards selected from a deck of 52. 1) How many different hands contain exactly 4 spades? (2) How many different hands contain exactly 4 cards of the same suit (any of the four suits)? (3) How many different hands contain cards of exactly one suit?

OpenStudy (ragingsquirrel):

You can only estimate this i believe...

OpenStudy (dumbcow):

1) 13C4 * 39C1 = 715 *39 = 27,885 2) 4 * 27,885 = 111,540 3) 4 * 13C5 = 4*1287 = 5,148